# Emulators for stellar profiles in binary population modeling

### v1

@@ 3.1.3 Post-processing
Some of the predicted profiles have some slight sinusoidal behavior as a result of the sinusoidal shapes of some of the principal components; we post-process the predicted profiles to enforce monotonic decreasing behavior so that the stellar density profile decreases from center to surface with no density inversions for all profiles. We accomplish this by starting at the surface and setting the value of any point whose value is lower than the previous point to that of the previous point. Since the profiles are renormalized according to interpolated central density, we also cut off any values above that value. An example of a predicted density profile before and after enforcing monotonic decreasing is shown in Figure 7. Another option for implementing this would be to start at the center and force rising profile points downward; as it turns out both options produce near-identical error distributions against both the unprocessed predictions and the true profiles.
@@ 4 Results
In this section we share quantitative information on predictive model evaluation. In order to show errors produced by the prediction methods presented here without the effect of errors from the scalar interpolators used for predicting quantities used in profile emulation, we use their true final central density values from the pre-computed test grid.
We also compare the error distributions of our profile predictions to the profile of the nearest neighbor binary in the initial parameter space. Like with the predictions, we rescale these nearest neighbor profiles to the true central density values to provide an even comparison. The other applications of neural network emulators in Ribas et al. (2020) and Auld et al. (2008) discussed in Section 3 do not provide such evaluations against nearest neighbor approximation. This is because these studies focus on calculating a posterior probability density and not on providing a surrogate for the original model. Therefore, the nearest neighbor approximation is not an appropriate tool for their problem and hence does not even represent a relevant benchmark. However, nearest neighbor is well-suited for population synthesis applications, so our predictions must compete with the accuracy of nearest neighbor profiles.
@@
$\eta_{i}=|\frac{\rho_{pred}(m_{enc,i})-\rho_{true}(m_{enc,i})}{\rho_{true}(m_{enc,i})}|.$ (4)
where $\rho_{pred}(m_{enc,i})$ and $\rho_{true}(m_{enc,i})$ are the predicted and true density values of the $i$-th profile point, respectively.
To measure the accuracy of an entire grid, we consider the distribution of all $\eta_{i}$ for all profiles within that grid. We also consider the median $\eta_{i}$ for each profile to be able to evaluate the accuracy of each profile. Both of these distributions are shown for each of the grids in Figures 8, 9, and 10, respectively.
@@
When it comes to comparing the machine learning model’s errors against nearest neighbor errors, we are most concerned with maximizing performance between 10% and 100% relative error. This is because improvements in performance below 10%, for example from 1% to 0.1%, are negligible given the errors introduced by the code. Changes in errors above 100%, for example between 200% and 2000%, are not as significant because 100% is already a very significant error. Therefore, it’s most important that a high percentage of profiles have relative error within 10%, and that the distribution between 10% and 100% skews more toward 10%.
@@
Figure 8: Cumulative density functions showing relative errors in HMS-HMS density profiles. We compare our predicted profiles in solid lines to rescaled nearest neighbor profiles in dotted lines. The distributions of all profile errors $\eta_{i}$ in the grid are in blue, whereas orange shows the distributions of the per-profile median $\eta_{i}$.
@@
Figure 8 shows that in the HMS-HMS grid, our best predictions are comparable to rescaled nearest neighbor profiles in the 10-100% range. This grid also has the least favorable performance of the three compared to nearest neighbor, potentially due to the higher level of complexity in binary evolution compared to CO-HMS and CO-HeMS.
@@
Figure 9: Cumulative density functions showing relative errors in CO-HMS density profiles. We compare our predicted profiles in solid lines to rescaled nearest neighbor profiles in dotted lines. The distributions of all profile errors $\eta_{i}$ in the grid are in blue, whereas orange shows the distributions of the per-profile median $\eta_{i}$.
@@
In Figure 9, we show that our predictions in the CO-HMS grid somewhat outperform rescaled nearest neighbor in the 10-100% relative error range.
@@
Figure 10: Cumulative density functions showing relative errors in CO-HeMS density profiles. We compare our predicted profiles in solid lines to rescaled nearest neighbor profiles in dotted lines. The distributions of all profile errors $\eta_{i}$ in the grid are in blue, whereas orange shows the distributions of the per-profile median $\eta_{i}$.
@@
We find that the CO-HeMS grid profiles are the least variable and complex out of the three grids, so there is the least room for improvement in the 10-100% relative error range, as shown in Figure 10. Our machine learning predictions for this grid have a marginal improvement at most in this range.
@@ 5 Conclusions
Being able to emulate final stellar profiles at the end of MESA evolutionary stages in POSYDON has a number of potential science applications. More information about the internal stellar profiles may provide insight into the behavior of binaries with anomalous properties. POSYDON users may also want to derive various physical quantities from emulated profiles instead of directly interpolating those quantities, particularly if a target property was not available in the original grid. They can also be used as an ingredient in custom physical prescriptions for defining the next steps of evolution, particularly in more uncertain phenomena such as common envelope evolution and supernovae. Rotation profiles in particular can also be leveraged to derive black hole spins of a population, which can be compared to observationally-derived distributions.
The PCA-neural network method applied to density profiles here is also well suited for other quantities with smooth profiles such as temperature. [^5] Angular rotation profiles, despite appearing relatively smooth, do not monotonically decrease or have any consistency in normalized shape like density and temperature and therefore are not well represented by PCA. Without normalizing the profile, the PCA focuses on shapes of stars with very high internal rotation and neglects the profiles of more slowly rotating stars. Such profiles may require more elaborate neural network architecture without dimensionality reduction. To improve emulation accuracy with this method, we could include more dimensions; our results with $n=8$ principal components are fairly successful without implementing any sophisticated model architecture. It would be possible to compensate for more principal components by optimizing the model further.
@@ Acknowledgements
This work is supported primarily by two sources: the Gordon and Betty Moore Foundation (PI Kalogera, grant awards GBMF8477 and GBMF12341) and the Swiss National Science Foundation (PI Tassos Fragos, project numbers PP00P2_211006 and CRSII5_213497). E.T, P.M.S, K.A.R., and M.S. were supported by the project numbers GBMF8477 and GBMF12341. Z.D. is grateful for support from the CIERA Board of Visitors Research Professorship. J.J.A. acknowledges support for Program number (JWST-AR-04369.001-A) provided through a grant from the STScI under NASA contract NAS5-03127. S.S.B., M.U.K., and Z.X. were supported by the project number PP00P2_211006. S.S.B. and M.M.B. were supported by the project number CRSII5_213497. M.M.B. is also supported by the Boninchi Foundation and the Swiss Government Excellence Scholarship. K.K. is supported by a fellowship program at the Institute of Space Sciences (ICE-CSIC) funded by the program Unidad de Excelencia María de Maeztu CEX2020-001058-M. Z.X. acknowledges support from the Chinese Scholarship Council (CSC). E.Z. acknowledges support from the Hellenic Foundation for Research and Innovation (H.F.R.I.) under the “3rd Call for H.F.R.I. Research Projects to support Post-Doctoral Researchers” (Project No: 7933). The computations were performed at Northwestern University on the Trident computer cluster (funded by the GBMF8477 award) and at the University of Geneva on the Yggdrasil computer cluster. This research was partly supported by the computational resources and staff contributions provided for the Quest high-performance computing facility at Northwestern University, jointly supported by the Office of the Provost, the Office for Research and Northwestern University Information Technology.

### v2

@@ 3.1.3 Post-processing
Some of the predicted profiles display slight sinusoidal behavior as a result of the sinusoidal shapes of the principal components; we post-process the predicted profiles to enforce monotonic decreasing behavior so that the stellar density profile decreases from center to surface with no density inversions for all profiles. We accomplish this by starting at the surface and setting the value of any point whose value is lower than the previous point to that of the previous point. Since the profiles are renormalized according to interpolated central density, we also cut off any values above that value. An example of a predicted density profile before and after enforcing monotonic decreasing is shown in Figure 7. Another option for implementing this would be to start at the center and force rising profile points downward; as it turns out both options produce near-identical error distributions against both the unprocessed predictions and the true profiles.
@@ 4 Results
In this section we share quantitative information on predictive model evaluation. To isolate the errors produced by the prediction methods presented in this work from those arising from the scalar interpolation methods published previously in Fragos et al. (2023), we use the true final central density values from the pre-computed test grid rather than interpolated values. This ensures that the errors evaluated in this study are solely attributable to the machine learning methods developed here, without influence from the scalar interpolation that would be used in practical profile emulation.
We also compare the error distributions of our profile predictions to the profile of the nearest neighbor binary in the initial parameter space. Like with the predictions, we rescale these nearest neighbor profiles to the true central density values to provide an even comparison. Many astrophysical applications of neural network emulators, such as Ribas et al. (2020) and Auld et al. (2008), do not provide such evaluations against nearest neighbor approximation because they focus on calculating a posterior probability density and not on providing a surrogate for the original model. However, as presented in Maltsev et al. (2024), nearest neighbor approximation is well-suited for population synthesis applications, so our emulator’s predictions must compete with the accuracy of nearest neighbor profiles.
@@
$\eta_{i}=\left|\frac{\rho_{pred}(m_{enc,i})-\rho_{true}(m_{enc,i})}{\rho_{true}(m_{enc,i})}\right|.$ (4)
where $\rho_{pred}(m_{enc,i})$ and $\rho_{true}(m_{enc,i})$ are the predicted and true density values of the $i$-th profile point, respectively. To measure the accuracy of the entire testing set, we consider the distribution of all $\eta_{i}$ for all profiles within the set. We also consider the median $\eta_{i}$ for each profile to be able to evaluate the accuracy of each profile. Both of these distributions are shown for each of the grids in Figures 8, 9, and 10, respectively. We use cumulative distribution functions (CDFs) to represent the error distributions. At each error value, the CDF is given by the fraction of errors in the distribution that fall below that value; essentially a histogram plotted cumulatively from lowest to highest error. When reading these plots to determine which method is more accurate, it is generally useful to look for which curve is farther to the left at a given percentile, signifying a lower error value at that percentile, or which curve is higher for a given error value, signifying that more errors in the distribution fall below that error value.
@@
When it comes to comparing the machine learning model’s errors against nearest neighbor errors, we are most concerned with maximizing performance between 10% and 100% relative error. This is because improvements in performance below 10%, for example from 1% to 0.1%, are negligible given the errors introduced by the code. When both the nearest neighbor profile and our predicted profile have errors below 10%, the marginal difference in accuracy between the two methods is not significant for our purposes. Changes in errors above 100%, for example between 200% and 2000%, are not as significant because 100% is already a very significant error. Likewise in this region, when errors from both methods exceed 100%, neither is usable, rendering any comparative advantage in accuracy irrelevant in such cases. So we focus on the 10% to 100% range, where the comparison is consequential for our purposes. This region is highlighted in yellow in Figures 8, 9, and 10. When evaluating the CDF plots through this lens, the method with the better errors in this area of interest will have a distribution curve that is higher than the other curve in the area between -1 and 0 on the x-axis.
@@
Figure 8: Cumulative density functions showing relative errors in HMS-HMS density profiles. We compare our predicted profiles in solid lines to rescaled nearest neighbor profiles in dotted lines. The distributions of all profile errors $\eta_{i}$ in the grid are in blue, whereas orange shows the distributions of the per-profile median $\eta_{i}$. The yellow highlighted area signifies the region of interest between 10% and 100% error.
@@
Figure 8 shows that in the HMS-HMS grid, our best predictions are comparable to rescaled nearest neighbor profiles in the highlighted 10-100% range. This grid also has the least favorable performance of the three compared to nearest neighbor, potentially due to the higher level of complexity in binary evolution compared to CO-HMS and CO-HeMS.
@@
Figure 9: Cumulative density functions showing relative errors in CO-HMS density profiles. We compare our predicted profiles in solid lines to rescaled nearest neighbor profiles in dotted lines. The distributions of all profile errors $\eta_{i}$ in the grid are in blue, whereas orange shows the distributions of the per-profile median $\eta_{i}$. The yellow highlighted area signifies the region of interest between 10% and 100% error.
@@
In Figure 9, we show that our predictions in the CO-HMS grid somewhat outperform rescaled nearest neighbor in the highlighted 10-100% relative error range.
@@
Figure 10: Cumulative density functions showing relative errors in CO-HeMS density profiles. We compare our predicted profiles in solid lines to rescaled nearest neighbor profiles in dotted lines. The distributions of all profile errors $\eta_{i}$ in the grid are in blue, whereas orange shows the distributions of the per-profile median $\eta_{i}$.The yellow highlighted area signifies the region of interest between 10% and 100% error.
@@
We find that the CO-HeMS grid profiles are the least variable and complex out of the three grids, so there is the least room for improvement in the highlighted 10-100% relative error range, as shown in Figure 10. Our machine learning predictions for this grid have a marginal improvement at most in this range.
@@ 5 Conclusions
Being able to emulate final stellar profiles at the end of MESA evolutionary stages in POSYDON has a number of potential science applications. Moreover, this study demonstrates the feasibility of a dimensionality reduction and neural network framework for stellar profile emulation in general. More information about the internal stellar profiles can provide insight into the behavior of binaries with anomalous properties. Being able to derive various physical quantities from emulated profiles instead of directly interpolating those quantities will expand the scope of binary population studies, particularly in cases where a target property was not saved in the original simulation. Emulated stellar profiles can also be used as an ingredient in custom physical prescriptions for defining the next steps of evolution, particularly in more uncertain phenomena such as common envelope evolution and supernovae. Rotation profiles in particular can also be leveraged to derive black hole spins of a population, which can be compared to observationally-derived distributions. By making our code publicly available, we aim to enable future studies to extend the approach presented here, paving the way for efficient and adaptable stellar profile emulators in diverse astrophysical contexts.
The PCA-neural network method applied to density profiles here combines principal component analysis (PCA) for dimensionality reduction with neural networks to make predictions in the lower-dimensional latent space defined by the principal components, enabling efficient modeling of high-dimensional data. This approach is also well suited for other quantities with smooth profiles such as temperature. [^5] Angular rotation profiles, despite appearing relatively smooth, do not monotonically decrease or have any consistency in normalized shape like density and temperature and therefore are not well represented by PCA. Without normalizing the profile, the PCA focuses on shapes of stars with very high internal rotation and neglects the profiles of more slowly rotating stars. Such profiles may require more elaborate neural network architecture without dimensionality reduction. To improve emulation accuracy with this method, we could include more dimensions; our results with $n=8$ principal components are fairly successful without implementing any sophisticated model architecture. It would be possible to compensate for more principal components by optimizing the model further.
@@ Acknowledgements
This work is supported primarily by two sources: the Gordon and Betty Moore Foundation (PI Kalogera, grant awards GBMF8477 and GBMF12341) and the Swiss National Science Foundation (PI Tassos Fragos, project numbers PP00P2_211006 and CRSII5_213497). E.T., P.M.S., K.A.R., and M.S. were supported by the project numbers GBMF8477 and GBMF12341. V.K. and A.K. gratefully acknowledge the support of the NSF-Simons AI-Institute for the Sky (SkAI) via grants NSF AST-2421845 and Simons Foundation MPS-AI-00010513. Z.D. is grateful for support from the CIERA Board of Visitors Research Professorship. J.J.A. acknowledges support for Program number (JWST-AR-04369.001-A) provided through a grant from the STScI under NASA contract NAS5-03127. S.S.B., M.U.K., and Z.X. were supported by the project number PP00P2_211006. S.S.B. and M.M.B. were supported by the project number CRSII5_213497. M.M.B. is also supported by the Boninchi Foundation and the Swiss Government Excellence Scholarship. K.K. is supported by a fellowship program at the Institute of Space Sciences (ICE-CSIC) funded by the program Unidad de Excelencia María de Maeztu CEX2020-001058-M. Z.X. acknowledges support from the Chinese Scholarship Council (CSC). E.Z. acknowledges support from the Hellenic Foundation for Research and Innovation (H.F.R.I.) under the “3rd Call for H.F.R.I. Research Projects to support Post-Doctoral Researchers” (Project No: 7933). The computations were performed at Northwestern University on the Trident computer cluster (funded by the GBMF8477 award) and at the University of Geneva on the Yggdrasil computer cluster. This research was partly supported by the computational resources and staff contributions provided for the Quest high-performance computing facility at Northwestern University, jointly supported by the Office of the Provost, the Office for Research and Northwestern University Information Technology.